# SoftShadow: Leveraging Penumbra-Aware Soft Masks for Shadow Removal

### v1

@@ Abstract
Recent advancements in deep learning have yielded promising results for the image shadow removal task. However, most existing methods rely on binary pre-generated shadow masks. The binary nature of such masks could potentially lead to artifacts near the boundary between shadow and non-shadow areas. In view of this, inspired by the physical model of shadow formation, we introduce novel soft shadow masks specifically designed for shadow removal. To achieve such soft masks, we propose a SoftShadow framework by leveraging the prior knowledge of pretrained SAM and integrating physical constraints. Specifically, we jointly tune the SAM and the subsequent shadow removal network using penumbra formation constraint loss and shadow removal loss. This framework enables accurate predictions of penumbra (partially shaded regions) and umbra (fully shaded regions) areas while simultaneously facilitating end-to-end shadow removal. Through extensive experiments on popular datasets, we found that our SoftShadow framework, which generates soft masks, can better restore boundary artifacts, achieve state-of-the-art performance, and demonstrate superior generalizability. 111 https://github.com/Xinrui014/SoftShadow/
@@ 1 Introduction
Shadow removal aims to restore content obscured in shadow regions and correct degraded illumination. Recently, deep learning methods have shown excellent performance in shadow removal tasks relying on large-scale training data. However, predicting shadow-free results remains challenging due to the highly variable illumination and color distortions in shadow areas. Boundary artifacts are particularly easy to occur because of the complex transitions in illumination at boundary regions, and it is also difficult to precisely locate the shadow areas around the boundary.
@@
Figure 1: Illustration of soft shadow removal results using our proposed SoftShadow with soft shadow mask compared to the most recent competing methods HomoFormer (Xiao et al. 2024) using hard shadow masks. The third column shows sharpened versions of our results and HomoFormer’s results for better visualization.
@@
A few methods rely on binary shadow masks as input to indicate the shadow regions. These binary shadow masks are either manually annotated (Wang, Li, and Yang 2018; Liu et al. 2024), or predicted by off-the-shelf shadow detectors (Cun, Pun, and Shi 2020). However, obtaining such shadow masks can be costly and complicated, making these methods less desirable for real-world applications. Moreover, binary masks fall short of representing the blurry edges of soft shadows, causing boundary artifacts in removal results. Soft shadows occur when the light source is partially occluded, resulting in a gradual transition between shadow and non-shadow regions, which is also known as the penumbra area. As shown in Figure 1, when the input image contains soft shadows, the previous state-of-the-art (SOTA) method (Xiao et al. 2024) suffers from boundary artifacts. Other approaches attempt shadow removal without explicitly extracting the shadow mask. Instead, they incorporate semantic parsing modules into shadow removal networks. For instance, some approaches involve predicting degradation attention (Qu et al. 2017; Cun, Pun, and Shi 2020), while others utilize domain classifiers (Jin, Sharma, and Tan 2021; Jin et al. 2024) to better understand and address shadow effects. However, their performance may be limited by the lack of powerful external detectors providing additional shadow location information, and the number of model parameters could increase significantly.
Figure 2: Illustration of the proposed SoftShadow. The left box illustrates the SoftShadow architecture, where a shadow image $\mathbf{y}$ is input into SAM for detecting soft shadow masks. The shadow removal network then processes the soft mask and the shadow image to produce a shadow-free image. The right box shows the three losses we used in SoftShadow. From top to bottom, the removal loss $\mathcal{L}_{rem}$ is calculated between shadow-free images and shadow-removal images. The mask loss $\mathcal{L}_{mask}$ is calculated between predicted soft masks and ground truth soft masks. The constraint loss $\mathcal{L}_{pen}$ is trying to regularize the gradient orientation and magnitude of predicted soft masks in the penumbra area.
@@
In view of this, we argue that simply leveraging the binary mask to represent the shadow location is not enough. Differently, we introduce a novel soft shadow mask specifically designed for shadow removal as shown in the hard mask v.s. soft mask of Figure 1, inspired by the physical shadow formation model. The soft masks can precisely locate the inner and outer boundaries of shadow and indicate the penumbra area with proper degradation variance. Moreover, recognizing the effectiveness of employing a powerful pre-trained detector for providing semantic information, we utilize the prior knowledge from the pre-trained SAM (Kirillov et al. 2023). While existing methods (Zhang, Gu, and Zhu 2023) have used the off-the-shelf SAM in a naive way, merely predicting the shadow mask as input to the shadow removal network, we adapt SAM as a soft shadow mask predictor and jointly optimize it with the shadow removal network.
@@
In this paper, we first propose a unified shadow removal framework, dubbed SoftShadow, which removes shadow from images in an end-to-end manner without requiring input shadow masks as shown in Figure 2. We integrate the pre-trained SAM combined with Low-Rank Adaptation (LoRA) (Hu et al. 2021) to jointly adapt the SAM model and shadow removal model at a much lower computational cost compared with the previous methods (Chen et al. 2023). Based on this, the framework enables accurate predictions of soft shadow masks with penumbra as intermediate results while simultaneously facilitating end-to-end shadow removal. Besides, we introduce a penumbra formation constraint to assist SAM adaptation. This constraint regularizes the gradient of the predicted mask in the penumbra area, resulting in more accurate predictions of detailed and spatially varied shadow masks. In the commonly used shadow removal datasets, SRD and LRSS datasets contain soft shadow images, which makes them more suitable for illustrating the improvements of utilizing soft shadow masks. Experimental results show that SoftShadow achieves superior performance on SRD and LRSS datasets. Our main contributions are summarized as follows:
@@
  • We propose a unified shadow removal framework that does not require additional input shadow masks, enabling accurate predictions of soft masks as intermediate results, which is specifically designed to capture detailed and spatially varied shadow location information.
@@
  • We introduce penumbra formation constraint inspired by physical shadow formation model to further refine the soft mask in the penumbra area. By leveraging the constraint loss and shadow removal loss, we jointly tune the SAM and the subsequent shadow removal network.
@@ 2.1 Shadow Removal
The degradation of shadows varies in each image, posing a significant challenge for shadow image restoration. In recent years deep learning-based approaches have achieved remarkable results in shadow removal (Guo et al. 2024). Some methods restore the shadow image with the guidance of shadow masks. For example, SP+M-Net (Le and Samaras 2019) employs two deep networks to predict shadow matte and shadow parameters. Recently, powerful backbones such as transformers (Vaswani et al. 2017; Dosovitskiy et al. 2020), and diffusion models (Ho, Jain, and Abbeel 2020; Saharia et al. 2022) have been applied to the shadow removal task. HomoFormer (Xiao et al. 2024) homogenizes the spatial distribution of shadow masks to uniformly recover the entire shadow image, while ShadowDiffusion (Guo et al. 2023b) provides a robust generative prior for producing natural shadow-free images. Some methods aim to eliminate the dependency on shadow mask inputs. DeShadowNet (Qu et al. 2017) introduces an end-to-end shadow removal method containing a multi-branch fusion module. Additionally, several methods utilize Generative Adversarial Networks (GANs) (Goodfellow et al. 2014) for shadow removal tasks. ST-CGAN (Wang, Li, and Yang 2018) connects two GANs in sequence to jointly detect and remove shadows. DC-ShadowNet (Jin, Sharma, and Tan 2021) provides an unsupervised domain-classifier discriminator for guided shadow removal network. More recently, DeS3 (Jin et al. 2024) implemented a method capable of removing shadows that are cast on the object itself without requiring shadow mask inputs. However, due to the absence of additional shadow location information, their performance may degrade, and the number of parameters used could increase significantly. Different from these existing methods, we introduce a new concept “soft mask”, designed explicitly for shadow removal.
Figure 3: (a) Illustration of the shadow formation geometry that creates the penumbra and umbra (Nielsen and Madsen 2007).; the penumbra area is where light is partially occluded. (b) Illustrate a shadow image example. (c) Illustrate the brightness curve at $H=175$. The penumbra area is clear in this curve.
@@ 2.2 Shadow Detection
Shadow detection can be approached as a segmentation task, it often struggles with small shadows and indistinct shadow edges. Various methods have been proposed to address these challenges. BDRAR (Zhu et al. 2018) introduces a bidirectional pyramidal architecture for shadow detection. DSD (Zheng et al. 2019) designs a distraction-aware module to minimize false positives in shadow detection. Chen et al. (Chen et al. 2020) propose a task-specific semi-supervised learning mechanism to utilize unlabeled shadow images for detection, enhancing the robustness of the model. DHAN (Cun, Pun, and Shi 2020) is a well-used shadow detection model using dual hierarchical aggregation network training on synthetic shadow and shadow-free image pairs to improve detection accuracy. The SAM-Adapter (Chen et al. 2023) has been utilized to fine-tune the Segment Anything Model (SAM) (Kirillov et al. 2023) specifically for shadow detection tasks. While this approach improves accuracy, it demands a substantial number of model parameters. However, most existing methods detect shadows as binary masks and fail to represent the penumbra area. In contrast, our approach leverages the shadow removal loss and the Penumbra Formation Constraint loss, jointly training the shadow removal network and the shadow detection network.
@@ 3 Methods
In this section, we first explain the motivation behind our newly introduced soft shadow mask, inspired by the physical shadow formation model. We then propose the unified shadow removal framework SoftShadow, which leverages the powerful segmentation capabilities of SAM to produce both the soft shadow mask and shadow-free images. Finally, we introduce the penumbra formation constraint loss to assist the SAM and incorporate physical constraints.
@@ 3.1 Motivation
Shadow degradations in the real world exhibit considerable variation and can be classified into two main categories: soft shadows and hard shadows (Langer and Zucker 1997). This classification depends on the light source and the distance between the object and the surface, as shown in Figure 3(a). Soft shadows are characterized by blurred edges and gradual transitions from light to dark, creating penumbra areas, while hard shadows have sharp edges without penumbra.
@@
Previous methods (Zhu et al. 2022; Fu et al. 2021; Guo et al. 2023a; Xiao et al. 2024) commonly use binary masks as guidance for shadow removal, effectively representing shadow locations with sharp edges. However, when applied to soft shadows with blurred boundaries, as shown in Figure 3 (b), binary masks struggle to fully capture these regions. The boundary between the shadow (umbra) and the non-shadow (lit) area is hard to determine. Additionally, we empirically observed that the decrease in brightness within the penumbra area (from lit to umbra area) is approximately linear, while the lit and umbra areas exhibit relatively uniform transitions, as shown in Figure 3(c).
@@
Based on the observations, we use soft mask $\mathbf{s}$ instead of binary mask $\mathbf{m}$. For a given shadow-free image $\mathbf{x}$, shadow image $\mathbf{y}$ is formulated as:
$\mathbf{y}=\mathbf{a}\cdot\mathbf{s}\cdot\mathbf{x}+(\mathbf{1}-\mathbf{s})\cdot\mathbf{x}$ (1)
@@
where $\cdot$ denotes element-wise multiplication, $\mathbf{s}$ is the soft shadow mask in which constant $s=0$ represents lit area, $s=1$ represents umbra area, and $s\in(0,1)$ represents penumbra area. $\mathbf{a}\in(0,1)$ represents the illumination weight in the non-lit area. The newly formulated shadow degradation model offers the potential for modeling soft masks more effectively.
3.2 Overall Architecture of SoftShadow
@@
SoftShadow is a unified framework designed to remove object shadows in images in an end-to-end manner. The overall architecture is illustrated in Figure 2. Within the SoftShadow framework, we leverage the segmentation capabilities of SAM by finetuning it to serve as our soft mask predictor. It was observed that the pretrained SAM often fails to produce object contours and accurately identify shadows (Jie and Zhang 2023). We use Low-Rank Adaptation (LoRA) (Hu et al. 2021) to finetune SAM with fewer trainable parameters and computational costs. We freeze all parameters in the image encoder and add a bypass on each self-attention block. Since the mask decoder is lightweight, we tune all the parameters of it.
@@
SAM is mainly trained for binary predictions, unlike the soft masks we aim to generate. To leverage SAM to generate continuous soft masks, we further introduce the penumbra formation constraints, i.e., using penumbra gradient loss $\mathcal{L}_{pen}$ and mask reconstruction loss $\mathcal{L}_{mask}$ to model the characteristics of soft masks in the penumbra area (details are described in Section 3.3). The constraints encourage the model to detect the shadow area and produce a continuous soft mask rather than a binary one, thereby capturing the shadow position more accurately.
Methods Input Masks shadow non-shadow all
[…]
SRD DHAN DHAN 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67
DC-ShadowNet NA 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65
BMNet DHAN 35.05 0.981 6.61 36.02 0.982 3.61 31.69 0.956 4.46
ShadowFormer DHAN 35.55 0.982 6.14 36.82 0.983 3.54 32.46 0.957 4.28
DeS3 NA 37.91 0.986 5.27 37.45 0.984 3.03 34.11 0.968 3.56
Homoformer DHAN 38.81 0.987 4.25 39.45 0.988 2.85 35.37 0.972 3.33
SAM-helps-shadow NA 33.94 0.979 7.44 33.85 0.981 3.74 30.72 0.952 4.79
ours NA 39.08 0.989 4.33 39.36 0.992 2.58 35.57 0.975 3.11
ISTD+ BMNet GT 37.87 0.991 5.62 37.51 0.985 2.45 33.98 0.972 2.97
ShadowDiffusion GT 39.69 0.992 4.97 38.89 0.987 2.28 35.67 0.975 2.72
HomoFormer GT 39.49 0.993 4.73 38.75 0.984 2.23 35.35 0.975 2.64
DC-ShadowNet NA 31.06 0.976 12.62 27.03 0.961 6.82 25.03 0.926 7.77
DeS3 NA 36.49 0.989 6.56 34.70 0.972 3.40 31.38 0.958 3.94
BMNet FDRNet - - 6.1 - - 2.9 - - 3.5
HomoFormer FDRNet 38.84 0.991 5.31 34.58 0.966 3.17 32.41 0.953 3.51
Ours NA 40.36 0.993 4.78 37.89 0.982 2.46 35.00 0.972 2.85
Table 1: The quantitative results of shadow removal using our SoftShadow and recent methods on SRD and ISTD+ datasets. The “Input Masks” column shows the different types of input masks used by these methods. “NA” means the method does not require masks as input. “GT” means the method uses manually annotated ground truth mask as input. The best and the second results are boldfaced and underlined, respectively.
Moreover, as shown in Figure 2, we feed the gradient of the shadow removal loss $\mathcal{L}_{rem}$ back to SAM to jointly train SAM and the subsequent shadow removal network:
$\mathcal{L}_{rem}=\mathbb{E}_{n}\left\|\mathbf{\hat{x}}_{n}-\mathbf{x}_{n}\right\|_{F}^{2}\;,$ (2)
where $n$ indexes all images, $\hat{\mathbf{x}}$ is the restored shadow-free image, and $\mathbf{x}$ is the ground truth shadow-free image. To this end, the intermediate soft shadow masks can be refined using shadow removal results, with the SAM adaptor optimized for better location guidance in subsequent shadow removal. The overall training objective of the proposed SoftShadow framework is:
where $\lambda_{1}$ and $\lambda_{2}$ are the weighting coefficients to balance the influence of each term. With the merits of joint training and physical constraints, our SAM adaptor (LoRA) can effectively predict soft masks.
3.3 Penumbra Formation Constraint
We introduce the Penumbra Formation Constraint to enable SAM to predict the soft masks more accurately. According to the shadow formation (Nielsen and Madsen 2007) as shown in Figure 3 (a), the light intensity in the penumbra area depends on the distance between the object and the surface where the shadow is cast. For simplicity, the light intensity $I(w)$ at point $w$ within the penumbra can be modeled by a linear function.
Figure 4: Illustrate the concept of Penumbra Formation Constraint. In the ideal soft shadow mask, the intensity in the penumbra area decreases progressively from the shadow center to the umbra area. The gradient orientation points from the shadow center towards the non-shadow area.
As shown in Figure 4, we assume that the direction from the shadow center to the shadow outward is positive. The light intensity is $I(w)=I_{0}\left(1-\frac{w_{0}-w}{L}\right)$, where $w$ is a pixel in the penumbra area, $w_{0}$ is the pixel at the edge between the penumbra and the lit area, and $L$ is the width of the penumbra area in the certain direction. $I_{0}$ is the light intensity in the non-shadow area. If set $I_{0}=1$, the soft mask intensity $I_{s}$ can be written as:
$I_{s}(w)=1-I(w)=\frac{w_{0}-w}{L}\;.$ (4)
The gradient of mask intensity is $\frac{dI_{s}}{dw}=-\frac{1}{L}$. The larger the absolute value of the gradient, the smaller the width $L$ of the penumbra region, which represents a smaller predicted penumbra area. Based on Eq (4), we introduce a penumbra gradient constraint to help the model predict the penumbra area more clearly. This constraint aims to ensure that the gradient direction in the penumbra area radiates outward from the shadow center and the gradient magnitudes are finite (when hard mask, the gradient magnitude is infinite). Since modeling the gradient in all directions from the shadow center is impractical. We calculate the mask gradient in the horizontal and vertical directions. The loss function can be formulated as:
$\mathcal{L}_{\text{pen}}=\mathbb{E}_{(n,p)}\left[R\left(o_{x}\frac{\partial I_{s_{n}}}{\partial x}\bigg{|}_{p}\right)+R\left(o_{y}\frac{\partial I_{s_{n}}}{\partial y}\bigg{|}_{p}\right)\right]\;,$ (5)
where n is the index of training images, and p is the pixel index in the penumbra area, $R$ is the ReLU function (Nair and Hinton 2010). For a predicted soft mask, we assume it nearly represents a complete object and identifies the shadow center $c$. Based on this center, we determine the orientation vector $o$ as:
$o_{x}=\begin{cases}1&s_{x}<c_{x},\\ -1&s_{x}\geq c_{x},\end{cases}\quad o_{y}=\begin{cases}1&s_{y}<c_{y},\\ -1&s_{y}\geq c_{y},\end{cases}$ (6)
In both Eq (5) and Eq (6), the variables $o$, $s$, and $c$ represent coordinate points, where the subscripts $x$ and $y$ indicate the values along the horizontal and vertical axes, respectively. $s$ represents the soft mask pixel in the penumbra area. We add a ReLU function to ignore pixels with values less than 00 and penalize the remaining values that are large, which indicates that the predicted intensity in the penumbra area changes quickly. We also employ a mask reconstruction loss $\mathcal{L}_{mask}$ to enable SAM to identify the approximate soft shadow contour as follows:

### v2

# SoftShadow: Leveraging Soft Masks for Penumbra-Aware Shadow Removal

@@ Abstract
Recent advancements in deep learning have yielded promising results for the image shadow removal task. However, most existing methods rely on binary pre-generated shadow masks. The binary nature of such masks could potentially lead to artifacts near the boundary between shadow and non-shadow areas. In view of this, inspired by the physical model of shadow formation, we introduce novel soft shadow masks specifically designed for shadow removal. To achieve such soft masks, we propose a SoftShadow framework by leveraging the prior knowledge of pretrained SAM and integrating physical constraints. Specifically, we jointly tune the SAM and the subsequent shadow removal network using penumbra formation constraint loss, mask reconstruction loss, and shadow removal loss. This framework enables accurate predictions of penumbra (partially shaded) and umbra (fully shaded) areas while simultaneously facilitating end-to-end shadow removal. Through extensive experiments on popular datasets, we found that our SoftShadow framework, which generates soft masks, can better restore boundary artifacts, achieve state-of-the-art performance, and demonstrate superior generalizability.
@@ 1 Introduction
Shadow removal aims to restore content obscured in shadow regions and correct degraded illumination. Recently, deep learning methods have shown excellent performance in shadow removal tasks relying on large-scale training data. However, one of the main challenges in shadow removal arises in images with soft shadows, where the shadow edges are blurred rather than sharp. This blurred area, known as the penumbra area, occurs due to partial occlusion of light, creating a transition zone where illumination varies significantly. Removing shadows in these regions is challenging because the gradual light-to-shadow transition often leads to boundary artifacts, compromising the quality of the shadow removal results.
@@
Figure 1: Illustration of soft shadow removal results using our proposed SoftShadow with soft shadow mask compared to the recent competing methods BMNet [40], SG-ShadowNet [31] and HomoFormer [35] using hard shadow masks. The second rows are sharpened versions of results for better visualization.
@@
A few methods rely on hard shadow masks (binary masks) to indicate the shadow regions [40, 8, 9, 35]. These binary shadow masks are either manually annotated [32, 23], or predicted by off-the-shelf shadow detectors [3]. Obtaining such shadow masks can be costly and complicated, and the choice of shadow detector can significantly impact shadow removal performance. Moreover, hard masks fall short of representing the penumbra area, causing boundary artifacts in removal results, especially when the penumbra areas are highly pronounced. As shown in Figure 1, when the input image contains penumbra regions, the previous state-of-the-art (SOTA) method [35] suffers from boundary artifacts. Other approaches attempt shadow removal without explicitly extracting the shadow mask. Instead, they incorporate semantic parsing modules into shadow removal networks. For instance, some approaches involve predicting degradation attention [27, 3], while others utilize domain classifiers [16, 17] to better understand and address shadow effects. However, their performance may be limited by the lack of powerful external detectors providing additional shadow location information.
@@
In view of this, we argue that simply leveraging the binary mask to represent the shadow location is not enough. Differently, we introduce a novel soft shadow mask (grayscale mask) specifically designed for shadow removal as shown in the hard mask v.s. soft mask in Figure 1. The soft masks can precisely locate the inner and outer boundaries of shadow and indicate the penumbra area with proper degradation variance. Moreover, recognizing the effectiveness of employing a powerful pretrained detector for providing semantic information, we utilize the prior knowledge from the pretrained SAM [19]. While existing methods [36] have used the off-the-shelf SAM in a naive way, merely predicting the shadow mask as input to the shadow removal network, we adapt SAM as a soft shadow mask predictor and jointly optimize it with the shadow removal network.
Figure 2: (a) Illustration of the shadow formation geometry that creates the penumbra and umbra [26] regions; the umbra area is where the light is fully occluded, the penumbra area is where light is partially occluded, the lit area is where light is not occluded. (b) Examples of soft shadow images and hard shadow images from commonly used datasets.
@@
In this paper, we first propose a unified shadow removal framework, dubbed SoftShadow, which removes shadow from images in an end-to-end manner without requiring input shadow masks as shown in Figure 3. We integrate the pretrained SAM [19] combined with Low-Rank Adaptation (LoRA) [13] to jointly adapt the SAM model and shadow removal model. The framework enables accurate predictions of soft shadow masks as intermediate results while simultaneously facilitating end-to-end shadow removal. Besides, we introduce a penumbra formation constraint to assist predict soft shadow masks. This constraint regularizes the gradient of the predicted mask in the penumbra area, resulting in more accurate predictions with detailed and spatially varied shadow masks. In the commonly used shadow removal datasets, SRD and LRSS datasets contain soft shadow images, which makes them more suitable for illustrating the improvements of utilizing soft shadow masks. Experimental results show that SoftShadow achieves superior performance on SRD and LRSS datasets. Our main contributions are summarized as follows:
@@
  • We propose a unified shadow removal framework that does not require additional input shadow masks. This framework enables accurate predictions of soft masks as intermediate results, which is specifically designed to capture detailed and spatially varied shadow location information.
@@
  • We introduce penumbra formation constraint inspired by the physical shadow formation model to further refine the soft mask in the penumbra area. By leveraging the constraint loss and shadow removal loss, we jointly tune the SAM and the subsequent shadow removal network.
@@ 2.1 Shadow Removal
The degradation of shadows varies in each image, posing a significant challenge for shadow image restoration. In recent years deep learning-based approaches have achieved remarkable results in shadow removal. Some methods restore the shadow image with the guidance of shadow masks. For example, SP+M-Net [21] employs two deep networks to predict shadow matte and shadow parameters. Recently, powerful backbones such as transformers [30, 4], and diffusion models [12, 28] have been applied to the shadow removal task. HomoFormer [35] homogenizes the spatial distribution of shadow masks to uniformly recover the entire shadow image, while ShadowDiffusion [9] provides a robust generative prior for producing natural shadow-free images. Some methods aim to eliminate the dependency on shadow mask inputs. DeShadowNet [27] introduces an end-to-end shadow removal method containing a multi-branch fusion module. ST-CGAN [32] connects two GANs [6] in sequence to jointly detect and remove shadows. DC-ShadowNet [16] provides an unsupervised domain-classifier discriminator for guided shadow removal network. More recently, DeS3 [17] implemented a method capable of removing shadows that are cast on the object itself without requiring shadow mask inputs. However, due to the absence of additional shadow location information, their performance may degrade, and the number of parameters used could increase significantly. Different from these existing methods, we introduce a new concept “soft mask”, designed explicitly for shadow removal.
Figure 3: Illustration of the proposed SoftShadow. The left box illustrates the SoftShadow networks, where a shadow image $\mathbf{y}$ is input into SAM for detecting soft shadow masks. The shadow removal network then processes the soft mask and the shadow image to produce a shadow-free image. The right box shows the three losses we used in SoftShadow. From top to bottom, the shadow removal loss $\mathcal{L}_{rem}$ is calculated between shadow-free images and shadow-removal images. The mask reconstruction loss $\mathcal{L}_{mask}$ is calculated between predicted soft masks and ours ground truth soft masks. The penumbra formation constraint loss $\mathcal{L}_{pen}$ act as a regularization term. It aims to regularize the gradient of predicted soft masks in the penumbra area. As shown in the soft mask intensity curve, the ideal mask intensity in the penumbra area should not be too large and the gradient orientation should be consistence.
@@ 2.2 Shadow Detection
Shadow detection can be approached as a segmentation task, it often struggles with small shadows and indistinct shadow edges. Various methods have been proposed to address these challenges. BDRAR [38] introduces a bidirectional pyramidal architecture for shadow detection. DSD [37] designs a distraction-aware module to minimize false positives in shadow detection. Chen et al. [2] propose a task-specific semi-supervised learning mechanism to utilize unlabeled shadow images for detection, enhancing the robustness of the model. DHAN [3] is a well-used shadow detection model using dual hierarchical aggregation network training on synthetic shadow and shadow-free image pairs to improve detection accuracy. The SAM-Adapter [1] has been utilized to fine-tune the Segment Anything Model (SAM) [19] specifically for shadow detection tasks. While this approach improves accuracy, it demands a substantial number of model parameters. However, most existing methods detect shadows as hard masks and fail to represent the penumbra area. In contrast, our approach leverages the shadow removal loss and the Penumbra Formation Constraint loss, jointly training the shadow removal network and the shadow detection network.
@@ 3 Methods
In this section, we first explain the motivation behind our newly introduced soft shadow mask, inspired by the physical model of shadow formation. We then propose the unified shadow removal framework SoftShadow, which leverages the powerful segmentation capabilities of SAM to produce the soft shadow mask and guide the removal network to generate better shadow-free images. Finally, to further mitigate boundary artifacts, we introduce the penumbra formation constraint loss which ensures a smoother transition between the shadow region and the non-shadow region, offering better guidance to the shadow removal network.
@@ 3.1 Motivation
Shadow degradations in the real world exhibit considerable amount of variation and can be classified into two main categories: soft shadows and hard shadows [20]. This classification depends on the light source and the distance between the object and the surface, as shown in Figure 2(a). Soft shadows are characterized by blurry edges and gradual transitions from light to dark, creating relatively large penumbra areas, while hard shadows have sharp edges with few to none penumbra regions.
@@
Previous methods [40, 5, 8, 35] commonly use hard masks to guide the shadow removal process. The hard masks effectively provide positional guidance for hard masks with sharp edges. However, when the penumbra regions are more pronounced, as shown in Figure 2 (b), the boundary between the umbra and the lit region is hard to determine, making hard masks unsuitable for representing such soft shadows. Additionally, we observed that, in general, the brightness transition within the penumbra area can vary significantly, whereas the brightness of the lit and the umbra areas are relatively uniform.
@@
To this end, we instead use soft (greyscale) masks to represent the position of the shadows. Formally, we denote the soft mask as $\mathbf{s}$. For a given shadow-free image $\mathbf{x}$, the shadow image $\mathbf{y}$ can be formulated as such:
$\mathbf{y}=\mathbf{a}\cdot\mathbf{s}\cdot\mathbf{x}+(\mathbf{1}-\mathbf{s})\cdot\mathbf{x},$ (1)
@@
where $\cdot$ denotes element-wise multiplication, $\mathbf{s}$ is the soft shadow mask in which constant $s=0$ represents lit area, $s=1$ represents umbra area, and $s\in[0,1]$ represents penumbra area. $a\in[0,1]$ represents the illumination weight in the non-lit area. Based on our experiment, we quantitatively and qualitatively found that the proposed shadow degradation model could lead to fewer boundary artifacts and generally better shadow removal results.
3.2 Architecture of SoftShadow
@@
SoftShadow is a unified framework designed to remove object shadows in images in an end-to-end manner. The overall architecture is illustrated in Figure 3. SoftShadow contains a soft shadow detector and a subsequent shadow removal network. The soft shadow detector leverages the strong segmentation capabilities of SAM and further finetuning it to serve as a soft mask detector. This is because the pretrained SAM can fail to accurately identify shadows [19, 15]. Moreover, since SAM has noticeably more parameters compared to conventional shadow detection networks, we use Low-Rank Adaptation (LoRA) [13] to reduce the number of trainable parameters. In practice, we tune all the self-attention blocks of image encoders using LoRA. Since the mask decoder is lightweight, we tune all the parameters of it.
@@
SAM is mainly trained for binary predictions, unlike the soft masks we aim to generate. To leverage SAM to generate continuous soft masks, we further introduce the mask reconstruction loss $\mathcal{L}_{mask}$ and the penumbra formation constraint loss $\mathcal{L}_{pen}$ to model the characteristics of soft masks (details are described in Section 3.3). They encourage the model to detect the shadow area and produce a continuous soft mask rather than a binary one, thereby capturing the shadow position more accurately. We formulate a mask reconstruction loss $\mathcal{L}_{mask}$ as follows: